# ActionPose: Pretraining 3D Human Pose Estimation with the Dark Knowledge of Action

### v1

@@ Abstract
2D-to-3D human pose lifting is an ill-posed problem due to depth ambiguity and occlusion. Existing methods relying on spatial and temporal consistency alone are insufficient to resolve these problems because they lack semantic information of the motions. To overcome this, we propose ActionPose, a framework that leverages action knowledge by aligning motion embeddings with text embeddings of fine-grained action labels. ActionPose operates in two stages: pretraining and fine-tuning. In the pretraining stage, the model learns to recognize actions and reconstruct 3D poses from masked and noisy 2D poses. During the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action labels. Additionally, our framework incorporates masked body parts and masked time windows in motion modeling to mitigate the effects of ambiguous boundaries between actions in both temporal and spatial domains. Experiments demonstrate the effectiveness of ActionPose, achieving state-of-the-art performance in 3D pose estimation on public datasets, including Human3.6M and MPI-INF-3DHP. Specifically, ActionPose achieves an MPJPE of 36.7mm on Human3.6M with detected 2D poses as input and 15.5mm on MPI-INF-3DHP with ground-truth 2D poses as input.
Introduction
@@
3D human pose estimation has been a significant research topic for better understanding human motion. It serves as a precursor for numerous downstream tasks such as human action recognition, human-computer interaction, and virtual reality. In monocular 3D human pose estimation, the input is derived from either 2D human poses or videos. However, converting 2D inputs to 3D outputs is an ill-posed problem due to depth ambiguity and occlusion, resulting in multiple possible 3D poses corresponding to a single 2D pose.
@@
To address this challenge, recent work leverages spatial and temporal consistency. Various architectures, including TCN-based (Pavllo et al. 2019; Cheng et al. 2020), GCN-based (Cai et al. 2019; Ci et al. 2019; Wang et al. 2020), and transformer-based models (Zhang et al. 2022; Zhu et al. 2023; Peng, Zhou, and Mok 2024), have been designed to capture this information. And pretraining strategies involving random frame masking, joint masking, and noise addition on large-scale dataset have been proposed to enhance learning of spatial and temporal information (Zhu et al. 2023).
Figure 1: (Left) A side view of a person clapping, with overlapping hand joints. (Right) A front view of a person performing ‘Cloud Hands’ with the hands about to cross.
However, there exists many scenarios where spatial and temporal information alone is insufficient. For example, when a person claps their hands, the hands may overlap as they move toward each other from a side view, making it difficult to determine which hand is closer to the viewer, as shown in Figure 1. Such a depth ambiguity can be resolved with the knowledge of the clapping motion. Similarly, in the ‘Cloud Hands’ form of Tai Chi, one hand might always appear in front of the other from a front view despite that it actually moves back and forth relative to the other. Knowledge of the ‘Cloud Hands’ semantic context helps resolve this ambiguity. Thus, compared to spatial and temporal consistency, action information provides richer semantic context.
One intuitive way to integrate action clues into 3D human pose estimation is through multi-task learning: training 3D human pose estimation and human action recognition tasks simultaneously. However, very few datasets are available with both pose and detailed action descriptions. Most human pose estimation datasets only contain high-level action labels (e.g., Human3.6M (Ionescu et al. 2014) and MPI-INF-3DHP (Mehta et al. 2017)) and lack critical semantic information. To the best of our knowledge, to date, datasets with more descriptive action labels, such as BABEL (Punnakkal et al. 2021) and KIT (Krebs et al. 2021), suffer from significant class imbalance. For example, the verbs ‘walk’ and ‘transition’ occur frequently while the frequencies of other classes are significantly less in BABEL (Punnakkal et al. 2021). Moreover, textual action descriptions are inherently non-unique. Describing the timing of movements for different body parts is also challenging, especially during concurrent actions, resulting in ambiguous temporal and spatial boundaries.
@@
This paper proposes a two-stage framework consisting of pretraining and fine-tuning to address the aforementioned challenges. During the pretraining stage, the model learns to recognize actions and reconstruct 3D poses from masked and noisy 2D poses. Rather than encoding action labels as one-hot vectors, as done in previous work (Luvizon, Picard, and Tabia 2018, 2021), we align the text embeddings of descriptive action labels with motion representations, leveraging the rich semantic information contained in actions. To tackle the issues of imbalanced data distribution in the pretraining dataset and the inherent uncertainty in human action recognition tasks, we replace the cross-entropy loss in infoNCE with a combination of KL-divergence and focal loss (Lin et al. 2017) to supervise multi-modal alignment. Additionally, inspired by MotionBERT (Zhu et al. 2023), we implement varying-size temporal window masking and body part-level masking during pretraining. This approach simulates scenarios where the spatial and temporal boundaries of an action are ambiguous and encourages the model to recognize the action even with partial information. By the end of the pretraining stage, the model is equipped to recognize actions. In the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action descriptions.
@@
  • To the best of our knowledge, this is the first work to perform multi-modal pretraining that directly aligns the text embedding of descriptive action labels with motion sequences for a 3D human pose estimation task.
@@
  • We propose a new pre-training and fine-tuning strategy for motion representation learning. By integrating multi-modal representation learning of motion with descriptive action labels and 3D human pose estimation, the model is trained to embed action cues into 3D pose estimation during the pretraining stage.
  • Experiments demonstrate the effectiveness of ActionPose, achieving state-of-the-art performance on public datasets. Specifically, ActionPose surpasses all existing methods, including those based on temporal information and diffusion-based approaches, on MPI-INF-3DHP. On Human3.6M, it significantly outperforms temporal information-based methods and achieves results comparable to diffusion-based methods with significantly lower inference overhead.
Related Work
3D Human Pose Estimation
@@
3D human pose estimation can be categorized into two main approaches: the first involves estimating 3D human poses directly from images (Li and Chan 2015; Sun et al. 2018; Moon, Chang, and Lee 2019), and the second involves lifting 2D poses to 3D poses (Martinez et al. 2017; Liu et al. 2020; Zhang et al. 2022; Peng, Zhou, and Mok 2024). This work falls into the second category. Within this category, various architectures have been proposed to capture spatial and temporal information in motion dynamics, including TCN-based (Pavllo et al. 2019; Cheng et al. 2020), GCN-based (Cai et al. 2019; Ci et al. 2019; Wang et al. 2020), and transformer-based (Zhang et al. 2022; Zhu et al. 2023; Peng, Zhou, and Mok 2024) models.
@@
Since action labels provide comprehensive summaries of motion, we adopt spatial and temporal transformers (Zhu et al. 2023) as our backbone for the pose encoder, which effectively learns and fuses motion representations across both temporal and spatial dimensions.
Skeleton-based Human Action Recognition
@@
Another task related to understanding human motion is human action recognition. Unlike 3D human pose estimation, which focuses on precise pose reconstruction, human action recognition aims to comprehend human motion at a higher level. Despite the different objectives, both tasks require an understanding of the spatial relationships between the movements of different body joints and the temporal dynamics within motion. Consequently, recent works have designed various architectures to capture the spatio-temporal information of motion dynamics for human action recognition including transformer-based (Ahn et al. 2023; Zhu et al. 2023; Do and Kim 2024), GCN-based (Yan, Xiong, and Lin 2018; Chen et al. 2021; Chi et al. 2022; Zhou et al. 2024), LSTM-based (Zhu et al. 2016; Liu et al. 2017), and CNN-based (Zhang et al. 2019; Xu et al. 2022).
Figure 2: The overall architecture of ActionPose.
Multi-Task Learning
@@
We use multi-task learning when the knowledge gained from one task can benefit other tasks (Zhang and Yang 2022). This is particularly relevant for human pose estimation and human action recognition. The work in (Luvizon, Picard, and Tabia 2018) was the first to highlight the interconnections between these tasks and attempted to estimate 2D/3D human poses alongside human action recognition. In their approach, both tasks share a common multi-task CNN, with human action recognition having an additional head to classify actions based on estimated poses and visual features. However, their method does not fully utilize the rich semantic information of actions, as the action classes are encoded as one-hot vectors. Similarly, (Luvizon, Picard, and Tabia 2021) aims to estimate 2D and 3D human poses and corresponding actions in real-time, but also encodes action classes as one-hot vectors.
Multi-Modal Representation Learning
@@
Multi-modal representation learning methods, such as CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021), have proven highly effective for tasks like image captioning, text-image retrieval, and zero-shot image classification. These methods work by aligning text embeddings from a text encoder with feature embeddings from an image encoder. This alignment allows the image encoder to capture and utilize rich semantic information from images, significantly enhancing its performance on downstream tasks.
Pioneer works point out that similar techniques could be applied to human motion related tasks, such as human motion generation (Tevet et al. 2022), human action recognition (Xiang et al. 2023) and human pose estimation (Zheng et al. 2023). Among them, our work is most related to GAP (Xiang et al. 2023) and ActionPrompt (Zheng et al. 2023).
To enable each body part to recognize actions in human action recognition tasks, GAP (Xiang et al. 2023) aligns features of individual body parts with the corresponding text descriptions. The global action information is then derived from a combination of all parts of body, functioning like a model zoo where each part serves as an action predictor. In contrast, our proposed method is simple yet effective. We train each part of the body to classify actions by randomly masking them during training, similar to dropout techniques in machine learning. While some body parts are randomly omitted during training, all parts are utilized during inference.
Another work that aligns human motion and text embedding for human pose estimation is ActionPrompt (Zheng et al. 2023). Their approach consists of two components: an action-related text-prompt block and an action-specific pose-prompt block. The action-related text-prompt block helps the pose encoder recognize corresponding action labels from Human3.6M dataset (Ionescu et al. 2014), while the action-specific pose-prompt block refines poses based on the predicted action labels. However, because the action labels in Human3.6M dataset are coarsely defined and the learned text templates are shared across all actions, the semantic information within their text embeddings is not as rich as ours, which utilize action labels from the motion-with-language dataset (Punnakkal et al. 2021). Additionally, ActionPrompt (Zheng et al. 2023) requires storing pose-prompt embeddings for all actions, which is not memory efficient when the number of actions is large.
Methodology
Network Architecture
@@
The overall objective of ActionPose is to enable 3D human pose estimation and text-motion alignment. The network consists of two encoders to extract features from text and pose data. Embedding pooling layers are then applied to generate global representations of the text and pose. These representations are used for text-motion alignment, and a regression head subsequently projects the pose features to estimate the 3D pose. The network architecture is illustrated in Figure 2.
Text and Pose Encoders
@@
The ActionPose network consists of two parallel BERT-style models (Devlin et al. 2018) operating over pose and text domains. The text stream utilizes three layers of transformer blocks, following the architecture of the original BERT (Devlin et al. 2018). The pose stream employs five layers of Spatial-Temporal transformer blocks, following MotionBERT (Zhu et al. 2023).
@@
Text Embedding and Pose Embedding Pooling Layers for Late Fusion
The purpose of these layers is to project the learned features of the pose class token <POSE> and the text class token <CLS> to the same feature space. These layers consist of multiple MLPs (Multi-Layer Perceptrons).
@@
Pose Estimation Regression Head
This structure is designed to project pose features $\{h_{p_{1}}^{l_{1}+l_{2}},h_{p_{2}}^{l_{1}+l_{2}},...,h_{p_{T}}^{l_{1}+l_{2}}\}$ to predict 3D pose sequence $\mathbf{X}\in\mathbb{R}^{T\times J\times 3}$.
Pretraining Tasks
We introduce two new pretraining strategies in addition to random joint and frame-level masked motion modeling: to address ambiguous temporal and spatial boundaries between actions and to improve the ability to classify actions using partial body poses.
@@
Motion Modeling with Masked Body Parts
We partition the human body into six parts: upper body, lower body, head, arms, legs, and hips, as shown in Figure 3. One part is then randomly chosen to be masked.
In addition to enabling each body part to contribute to action classification, as discussed in Related WorkMulti-Modal Representation Learning and to mitigate the problem of ambiguous spatial boundaries between concurrent actions, this technique also helps the pose encoder learn the kinematics and anatomical details of the human body more comprehensively than joint-level masking alone, capturing the coordinated movements of multiple joints. When masked body part motion modeling is combined with multi-modal alignment prediction, the pose encoder further learns the spatial relationships of human motion within the context of corresponding actions. This masking technique also simulates real-world scenarios where parts of the human body may be occluded in a camera’s field of view.
Motion Modeling with Masked Time Windows
In addition to masked body parts, we propose another masking technique that masks the motion in the temporal domain with a varying window size between $[T_{1},T_{2}]$. The starting frame of this masked time window also varies and is randomly chosen.
@@
Multi-modal Alignment Prediction
The dual-stream encoders are optimized together by contrasting the global representation of text embeddings and pose embeddings at sample $j$:
@@
Since motion-and-language datasets (Punnakkal et al. 2021; Krebs et al. 2021) are highly imbalanced and human action recognition tasks are inherently non-deterministic, namely, motion and action labels do not have a one-to-one correspondence, we replace the cross-entropy in the infoNCE loss with a combination of focal loss (Lin et al. 2017) and KL divergence. The modified loss function is as follows:
@@
Objective Function
During pretraining, we adopt several masking strategies: joint-level and frame-level masking 50% of the time to capture spatial consistency among neighboring joints and temporal consistency across adjacent frames; random body part masking 25% of the time; and random time window masking 25% of the time. Additionally, we introduce noises into the 2D pose sequences.
@@
$\mathcal{L}_{3D}=\sum_{k=1}^{M}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{\hat{X}_{t,j}^{k}}-\mathbf{X_{t,j}^{k}}||_{2},$ (3)
$\mathcal{L}_{v}=\sum_{k=1}^{M}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{\hat{V}_{t,j}^{k}}-\mathbf{V_{t,j}^{k}}||_{2},$ (4)
@@
$\mathcal{L}=\mathcal{L}_{con}+\lambda_{3D}\mathcal{L}_{3D}+\lambda_{v}\mathcal{L}_{v}.$ (5)
Fine-tuning
During pretraining, the parameters of both the pose encoder and text encoder are updated, but during fine-tuning, only the pose encoder is fine-tuned on the target dataset. Specifically, 2D poses are detected from videos and then lifted to 3D poses. The fine-tuning process is supervised by the reconstruction losses as specified in Equation 3 and Equation 4.
Method MPJPE P-MPJPE
P-STMO (Shan et al. 2022) 42.8 34.4
MixSTE (Zhang et al. 2022) 40.9 32.6
GLA-GCN (Yu et al. 2023) 44.4 34.8
STCFormer (Tang et al. 2023) 40.5 31.8
MotionBERT (Zhu et al. 2023$\diamond$ 37.5
DiffPose (Gong et al. 2023) †* 36.9
KTPFormer (Peng, Zhou, and Mok 2024) 40.1 31.9
KTPFormer (Peng, Zhou, and Mok 2024) †* 33.0 26.2
ActionPose (Ours) $\diamond$ 36.7 31.3
Table 1: Quantitative comparison of 3D human pose estimation on the Human3.6M dataset using detected 2D poses. Errors are reported as average MPJPE (mm) and P-MPJPE (mm). All methods use a temporal window of 243 frames. †indicates the use of temporal information, * represents diffusion-based methods, and $\diamond$ indicates pretraining-based methods. Results from other models are directly quoted from the respective papers.
Method PCK $\uparrow$ AUC $\uparrow$ MPJPE $\downarrow$
P-STMO (Shan et al. 2022) 97.9 75.8 32.2
GLA-GCN (Yu et al. 2023) 98.5 79.1 27.7
STCFormer (Tang et al. 2023) 98.7 83.9 23.1
DiffPose (Gong et al. 2023) †* 98.0 75.9 29.1
KTPFormer (Peng, Zhou, and Mok 2024) †* 98.9 85.9 16.7
MotionAGFormer-L (Mehraban, Adeli, and Taati 2024) 98.2 85.3 16.2
ActionPose (Ours) $\diamond$ 98.9 87.0 15.5

### v2

# LangPose: Language-Aligned Motion for Robust 3D Human Pose Estimation

@@ Abstract
2D-to-3D human pose lifting is an ill-posed problem due to depth ambiguity and occlusion. Existing methods relying on spatial and temporal consistency alone are insufficient to resolve these problems especially in the presence of significant occlusions or high dynamic actions. Semantic information, however, offers a complementary signal that can help disambiguate such cases. To this end, we propose LangPose, a framework that leverages action knowledge by aligning motion embeddings with text embeddings of fine-grained action labels. LangPose operates in two stages: pretraining and fine-tuning. In the pretraining stage, the model simultaneously learns to recognize actions and reconstruct 3D poses from masked and noisy 2D poses. During the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action labels. Additionally, our framework incorporates masked body parts and masked time windows in motion modeling, encouraging the model to leverage semantic information when spatial and temporal consistency is unreliable. Experiments demonstrate the effectiveness of LangPose, achieving SOTA level performance in 3D pose estimation on public datasets, including Human3.6M and MPI-INF-3DHP. Specifically, LangPose achieves an MPJPE of 36.7mm on Human3.6M with detected 2D poses as input and 15.5mm on MPI-INF-3DHP with ground-truth 2D poses as input.
1 Introduction
@@
3D human pose estimation (HPE) is crucial for better understanding human motion and serves as a cornerstone for numerous downstream tasks such as human action recognition, human-computer interaction, and virtual reality. 2D-to-3D pose lifting, which estimates 3D human poses from 2D keypoints extracted from a single image or video frame, is a mainstream approach to 3D human pose estimation due to the maturity and computational efficiency of modern 2D pose detectors. However, it is an ill-posed problem as depth ambiguity and occlusion often result in multiple possible 3D poses corresponding to a single 2D pose.
@@
To cope with the inherent ambiguity, prior work has relied on spatial and temporal cues to enforce consistency in 3D pose predictions. Various architectures, including TCN-based [35, 4], GCN-based [2, 6, 49], and transformer-based models [55, 62, 36], have been designed to capture this information. Pretraining strategies such as random frame masking, joint masking, and noise addition on large-scale dataset have been utilized to enhance the learning of spatial and temporal cues [62].
However, spatial and temporal consistency diminishes in the presence of significant occlusions or high dynamic actions such as those found in the real world (e.g., in sports). Semantic information refers to high-level contextual knowledge about human actions, intentions, or interactions. Such information, often available in fine-grained textual descriptions of human activities, can inform or constrain pose estimation and help disambiguate pose configurations, particularly in cases of occlusion and depth ambiguity. Semantic information can serve as a complementary cue. For example, when only the upper body is visible, semantic priors associated with the action “sitting down” indicate that the hips are flexed and the knees are bent to roughly $90^{o}$, allowing the model to plausibly reconstruct the occluded lower‑limb joints. Similarly, in highly dynamic scenes, where temporal cues become less reliable, semantic priors linked to the ongoing action, for example, the expected limb trajectories and contact events in “dribbling a ball”, can guide the estimation of a full pose sequence even if it is partially occluded over time.
One possible approach to incorporating action cues into 3D human pose estimation is through multi-task learning (MTL), where pose estimation and action recognition are jointly optimized. However, MTL typically represents action categories with one‑hot vectors, thereby discarding the rich semantic relationships conveyed in fine‑grained textual descriptions. Furthermore, such models do not generalize well to pose sequences from unseen action classes. Therefore, we argue for explicitly aligning the latency feature space of pose sequences with action descriptions and propose LangPose, a framework that learns a shared pose–language embedding via cross-modal contrastive training. Datasets designed for human motion generation, such as BABEL [40] and KIT [18], offer richer and more fine-grained action descriptions. However, they introduce new challenges: (1) severe class imbalance—word like ‘walk’ and ‘transition’ occur with high frequency, while many other descriptions are underrepresented in BABEL [40]; and (2) the non-uniqueness of textual action descriptions, as the same motion can correspond to multiple valid descriptions, especially in the presence of concurrent or compound actions, and vice versa.
@@
To address the above challenges, LangPose leverages a two-stage training process comprising a pretraining phase followed by fine-tuning. In the pretraining stage, the model is trained on a synthetic dataset [40] containing motion sequences paired with fine-grained action labels, jointly optimizing for semantic alignment and 3D pose reconstruction. The semantic knowledge acquired during pretraining is generalizable across different 3D pose estimation datasets, in contrast to previous methods [26, 27, 59] that require retraining an action classifier for each target dataset. To tackle the issues of imbalanced data distribution in the pretraining dataset and the inherent uncertainty in human action recognition tasks, we replace the cross-entropy loss in infoNCE with a combination of KL-divergence and focal loss [22] to supervise multi-modal alignment. Additionally, inspired by MotionBERT [62], we implement varying-size temporal window masking and body part-level masking during pretraining. These strategies further encourage the model to reconstruct poses given semantic information rather than relying solely on spatial and temporal consistency. By the end of the pretraining stage, the model is equipped to recognize actions. In the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action descriptions.
@@
  • To the best of our knowledge, this is the first work to perform multi-modal pretraining that directly aligns the text embedding of descriptive action labels with skeleton motion sequences for a 3D human pose estimation task.
@@
  • We propose a new pre-training and fine-tuning strategy for motion representation learning. By integrating multi-modal representation learning of motion with descriptive action labels and 3D human pose estimation, the model is trained to embed generalizable action cues into 3D pose estimation during the pretraining stage.
  • Experiments show that LangPose achieves competitive results on public benchmarks, surpassing most methods on MPI-INF-3DHP and attaining SOTA-level performance on Human3.6M, with notable gains on challenging actions and lower inference cost than diffusion-based models.
2 Related work
2.1 3D Human pose estimation
@@ 2 Related work
3D human pose estimation can be categorized into two main approaches: the first involves estimating 3D human poses directly from images [19, 45, 33], and the second involves lifting 2D poses to 3D poses [29, 25, 55, 36]. This work falls into the second category. Within this category, various architectures have been proposed to capture spatial and temporal information in motion dynamics, including TCN-based [35, 4], GCN-based [2, 6, 49], and transformer-based [55, 62, 36, 24] models.
@@
Since action labels provide comprehensive summaries of motion, we adopt spatial and temporal transformers [62] as our backbone for the pose encoder, which effectively learns and fuses motion representations across both temporal and spatial dimensions.
2.2 Skeleton-based human action recognition
@@
Another task related to understanding human motion is human action recognition. Unlike 3D human pose estimation, which focuses on precise pose reconstruction, human action recognition aims to comprehend human motion at a higher level. Despite the different objectives, both tasks require an understanding of the spatial relationships between the movements of different body joints and the temporal dynamics within motion. Consequently, recent works have designed various architectures to capture the spatio-temporal information of motion dynamics for human action recognition including transformer-based [1, 62, 10], GCN-based [53, 3, 5, 60], LSTM-based [61, 23], and CNN-based [56, 52].
Figure 1: The overall architecture of LangPose.
2.3 Multi-task learning
@@
We use multi-task learning when the knowledge gained from one task can benefit other tasks [57]. This is particularly relevant for human pose estimation and human action recognition. The work in [26] was the first to highlight the interconnections between these tasks and attempted to estimate 2D/3D human poses alongside human action recognition. In their approach, both tasks share a common multi-task CNN, with human action recognition having an additional head to classify actions based on estimated poses and visual features. However, their method does not fully utilize the rich semantic information of actions, as the action classes are encoded as one-hot vectors. Similarly, [27] aims to estimate 2D and 3D human poses and corresponding actions in real-time, but also encodes action classes as one-hot vectors.
2.4 Multi-modal representation learning
@@
Multi-modal representation learning methods, such as CLIP [41] and ALIGN [16], have proven highly effective for tasks like image captioning, text-image retrieval, and zero-shot image classification. These methods work by aligning text embeddings from a text encoder with feature embeddings from an image encoder. This alignment allows the image encoder to capture and utilize rich semantic information from images, significantly enhancing its performance on downstream tasks.
Pioneer works point out that similar techniques can be applied broadly across the motion–language spectrum, including text-to-motion generation [48, 38, 37], motion captioning [11], text–motion retrieval [39, 32], and text-aided discriminative tasks such as action recognition [51] and pose estimation [59, 8, 12]. Building on these task-specific advances, a recent trend is to pursue general-purpose motion–language models that unify generation, captioning, retrieval, and discriminative perception within a single pretrained backbone (e.g., shared encoders with lightweight task heads) [20, 17, 50]. In parallel, substantial effort has gone into curating large-scale motion–language corpora and whole-body resources to enable such unified modeling [14, 18, 40, 21, 58].
Among the studies above, the most relevant to our setting are two lines of work that reconstruct 3D motion from different inputs: multi-modal 3D human pose estimation from 2D to 3D, and text-conditioned motion generation from text to 3D. Our approach differs in two ways. First, prior multi-modal 3D HPE with textual cues usually centers on images rather than videos [8, 12], so the semantics of motion dynamics remain underexplored for the 3D HPE task. Second, our method complements text-conditioned generation without performing generation: rather than synthesizing motion from language, we leverage motion semantics captured by text–motion alignment as auxiliary supervision to regularize 2D-to-3D pose estimation, preserving camera-aware, metric accuracy while adding semantic guidance, especially under occlusion. In this way, we complete the picture—from learning semantics to using semantics for reconstruction.
The most closely related approach that aligns motion and text for human pose estimation is ActionPrompt [59]. Their method consists of two main components: an action-related text-prompt block and an action-specific pose-prompt block. The former enables the pose encoder to predict action labels from pose sequences using supervision from human pose estimation datasets [15, 44], while the latter refines the predicted poses based on the identified action. However, the semantic representations learned by the action-related text-prompt block are not transferable across datasets and must be retrained for each new domain. Moreover, since action labels in pose estimation datasets are typically coarse, this approach relies heavily on the action-specific pose prompts to recover fine-grained semantic details. This two-stage pipeline further assumes accurate classification of motion into discrete action labels, which is prone to failure in the presence of severe noise. In contrast, LangPose performs end-to-end alignment between motion and semantic information while reconstructing 3D pose sequences. The learned semantics are generalizable across datasets and eliminate the need for storing action-specific pose prompts during inference.
3 Methodology
3.1 Network architecture
@@
The overall objective of LangPose is to enable 3D human pose estimation and text-motion alignment. The network consists of two encoders to extract features from text and pose data. Embedding pooling layers are then applied to generate global representations of the text and pose. These representations are used for text-motion alignment, and a regression head subsequently projects the pose features to estimate the 3D pose. The network architecture is illustrated in Fig. 1.
@@
Text and pose encoders. The LangPose network consists of two parallel BERT-style models [9] operating over pose and text domains. The text stream utilizes three layers of transformer blocks, following the architecture of the original BERT [9]. The pose stream employs five layers of Spatial-Temporal transformer blocks, following MotionBERT [62].
@@
Text embedding and pose embedding pooling layers for late fusion. The purpose of these layers is to project the learned features of the pose class token <POSE> and the text class token <CLS> to the same feature space. These layers consist of multiple MLPs (Multi-Layer Perceptrons).
@@
Pose estimation regression head. This structure is designed to project pose features $\{h_{p_{1}}^{l_{1}+l_{2}},h_{p_{2}}^{l_{1}+l_{2}},...,h_{p_{T}}^{l_{1}+l_{2}}\}$ to predict 3D pose sequence $\mathbf{X}\in\mathbb{R}^{T\times J\times 3}$.
3.2 Pretraining tasks
We introduce two new pretraining strategies in addition to random joint and frame-level masked motion modeling: to promote semantic reasoning under weak spatial and temporal consistency and to improve the ability to classify actions using partial body poses.
@@ 3.2 Pretraining tasks
Motion modeling with masked body parts. We partition the human body into six parts: upper body, lower body, head, arms, legs, and hips, as shown in Fig. 2. One part is then randomly chosen to be masked.
This technique enables the pose encoder to learn human kinematics and anatomy more comprehensively than joint-level masking by capturing coordination not only among individual joints but also across multiple body parts. When masked body part motion modeling is combined with multi-modal alignment prediction, the pose encoder further learns the spatial relationships of human motion within the context of corresponding actions. This masking technique also simulates real-world scenarios where parts of the human body may be occluded in a camera’s field of view.
Motion modeling with masked time windows. In addition to masked body parts, we propose another masking technique that masks the motion in the temporal domain with a varying window size between $T_{1},T_{2}]$. The starting frame of this masked time window also varies and is randomly chosen.
@@
Multi-modal alignment prediction. The dual-stream encoders are optimized together by contrasting the global representation of text embeddings and pose embeddings at sample $j$:
@@
Since motion-and-language datasets [40, 18] are highly imbalanced and human action recognition tasks are inherently non-deterministic, namely, motion and action labels do not have a one-to-one correspondence, we replace the cross-entropy in the infoNCE loss with a combination of focal loss [22] and KL divergence. The modified loss function is as follows:
@@
Objective function. During pretraining, we adopt several masking strategies: joint-level and frame-level masking 50% of the time to capture spatial consistency among neighboring joints and temporal consistency across adjacent frames; random body part masking 25% of the time; and random time window masking 25% of the time. Additionally, we introduce noises into the 2D pose sequences.
@@
[…]
@@
$\mathcal{L}=\lambda_{con}\mathcal{L}_{con}+\lambda_{3D}\mathcal{L}_{3D}+\lambda_{v}\mathcal{L}_{v}.$ (5)
3.3 Fine-tuning
During pretraining, the parameters of both the pose encoder and text encoder are updated, but during fine-tuning, only the pose encoder is fine-tuned on the target dataset. Specifically, 2D poses are detected from videos and then lifted to 3D poses. The fine-tuning process is supervised by the reconstruction losses as specified in Eq. 3 and Eq. 4.
Table 1: Quantitative comparison of 3D human pose estimation on the Human3.6M dataset using detected 2D poses. The per-action performance is shown under P1, and the average metrics across P1, P2, and P3 are summarized. All methods use a temporal window of 243 frames. †indicates the use of temporal information, * represents diffusion-based methods, and $\diamond$ indicates pretraining-based methods. Results from other models are directly quoted from the respective papers.
Method Dir. Disc. Eat Greet Phone Photo Pose Pur. Sit SitD. Smoke Wait WalkD. Walk WalkT. P1 ↓ P2 ↓ P3 ↓
P-STMO [42] 38.9 42.7 40.4 41.1 45.6 49.7 40.9 39.9 55.5 59.4 44.9 42.2 42.7 29.4 29.4 42.8 34.4 54.9
MixSTE [55] 37.6 40.9 37.3 39.7 42.3 49.9 40.1 39.8 51.7 55.0 42.1 39.8 41.0 27.9 27.9 40.9 32.6 52.2
GLA-GCN [54] 41.3 44.3 40.8 41.8 45.9 54.1 42.1 41.5 57.8 62.9 45.0 42.8 45.9 29.4 29.9 44.4 34.8 58.3
STCFormer [47] 38.4 41.2 36.8 38.0 42.7 50.5 38.7 38.2 52.5 56.8 41.8 38.4 40.2 26.2 27.7 40.5 31.8 53.3
KTPFormer [36] 37.3 39.2 35.9 37.6 42.5 48.2 38.6 39.0 51.4 55.9 41.6 39.0 40.0 27.0 27.4 40.1 31.9 51.8
TCPFormer [24] 37.9 31.7
RePose [46] 34.6 36.8 35.0 31.3 38.8 44.7 35.3 35.5 47.5 50.9 38.4 35.5 34.9 23.9 25.0 36.5 30.7 47.7
DiffPose [13] †* 33.2 36.6 33.0 35.6 37.6 45.1 35.7 35.5 46.4 49.9 37.3 35.6 36.5 24.4 24.1 36.9 47.1
D3DP [43] †* 37.3 39.5 35.6 37.8 41.3 48.2 39.1 37.6 49.9 52.8 41.2 39.2 39.4 27.2 27.1 39.5 31.6 50.3
ActionPrompt [59] 37.7 40.2 39.8 40.6 43.1 48.0 38.8 38.9 50.8 63.2 42.0 40.0 42.0 30.5 31.6 41.8 54.0
[…]